IRON MODIFICATION OF TiO_{2}

\underline{NKENG ASONG}, FAITH DUKES, MARY JANE SHULTZ, Department of Chemistry, Tufts University, Medford, MA 02155.

TiO_{2} is a photocatalyst that has many applications particularly in the environmental field. TiO_{2} and doped TiO_{2} (0.1%, 0.5%, 1%, 2.5% and 5% Fe) with particle sizes ranging from 2-4 nm are synthesized by the inorganic precursor TiCl_{4}. Results from electron paramagnetic resonance suggest that the distribution of iron is on the surface of the doped crystals. Sum frequency generation is used to probe methanol adsorption on these catalysts. The SFG spectra obtained show that the methoxy peak in undoped TiO_{2} is suppressed in the doped crystals. These results have great implications in the tailoring of the photocatalyst.